OpenStudy (anonymous):

**will give medal** Write the point-slope form of the equation passing through (5, -1) with a slope of 6. A. y-1=6(x+5) B. y+5=6(x-1) C. y+1=6(x-5) D. y-5=6(x+1)

jimthompson5910 (jim_thompson5910):

Point slope form is \[\Large y-y_{1} = m(x-x_{1})\] Plug in the given slope m = 6 and the given point \(\Large (x_{1},y_{1}) = (5,-1)\) to get \[\Large y-(-1) = 6(x-5)\] \[\Large y+1 = 6(x-5)\]